OpenStudy (anonymous):

the lins passes through points (x,4) and (4,-5) is perpindiculr to to a line with a slope of -7/3 what is the value of x ?

OpenStudy (kropot72):

I just replied to this same problem in closed questions.

OpenStudy (kropot72):

The slope of the line with the given points is 3/7 the reason being the line is perpendicular to the line with slope -7\3 \[4-(-5)=\frac{3}{7}(x-4)\] \[9+\frac{12}{7}=\frac{3x}{7}\] \[\frac{7}{3}(9+\frac{12}{7})=x\] 21 + 4 = 25 = x